Calculate the magnitude of the electric field 2.0 m from a long wire that is charged uniformly at λ 106 C/m. 4.0 x E = 6.3 × 104 N/C E 3.6 X 101 N/C E-2.6 × 104 N/C U/ E 4.9 )x 104 N/C Question 6 (12.5 points) Two charges ± 10 pC are places placed 5.00 mm apart. Determine the magnitude of the electric field at a point P on the axis of the dipole. 15.0 cm away from the center of the dipole. E= 7.2× 105 N/C E = 2.7× 105 N/C E-5.4x 105 N/C E 1.4X 103 N/C
Answer #1

Solution:

a) Using the formula:

2TEQT 07

=> E = 4 * 10^-6 / (2 * 3.14 * 8.84 * 10^-12 * 2)

=> E = 3.6 * 10^4 N/C ( approx)

b) Using the formula,

=> E = K * 2P / r^3

=> E = 9 * 10^8 * 2 * 2 * 10^-5 * 5 * 10^-3 / (15*10^-2)^3)

=> E = 2.7 * 10^5 N/C

please comment for queries.

thanks.